# Leveraging AI to Accelerate Clinical Data Cleaning: A Comparative Study of AI-Assisted vs. Traditional Methods

### v1

@@ Abstract.
Clinical trial data cleaning represents a critical bottleneck in drug development, with manual review processes struggling to manage exponentially increasing data volumes and complexity. This paper presents Octozi, an artificial intelligence-assisted platform that combines large language models with domain-specific heuristics to transform clinical data review. In a controlled experimental study with experienced clinical reviewers (n=10), we demonstrate that AI assistance increased data cleaning throughput by 6.03-fold while simultaneously decreasing cleaning errors from 54.67% to 8.48% (a 6.44-fold improvement). Crucially, the system reduced false positive queries by 15.48-fold, minimizing unnecessary site burden. These improvements were consistent across reviewers regardless of experience level, suggesting broad applicability. Our findings indicate that AI-assisted approaches can address fundamental inefficiencies in clinical trial operations, potentially accelerating drug development timelines and reducing costs while maintaining regulatory compliance. This work establishes a framework for integrating AI into safety-critical clinical workflows and demonstrates the transformative potential of human-AI collaboration in pharmaceutical clinical trials.
@@ 1. Introduction
The pharmaceutical industry faces an unprecedented data management crisis. Clinical trials now generate three times more data than a decade ago (Taves, 2010), yet the fundamental processes for ensuring data quality remain largely unchanged since the advent of electronic data capture (EDC). This disconnect between data volume and processing capability creates substantial delays in drug development, with each day of database lock delay potentially costing \$500,000 in lost revenue across a product’s lifecycle (Smith et al., 2024). More critically, inadequate data cleaning risks missing crucial safety signals that could impact patient welfare.
Clinical data cleaning, the systematic process of reviewing, validating, and reconciling data collected from investigative sites, serves as the foundation for regulatory submissions to authorities including the US Food and Drug Administration (FDA) and European Medicines Agency (EMA). This process extends beyond simple error detection to encompass complex medical judgment, requiring specialized expertise to identify clinically meaningful discrepancies while maintaining regulatory compliance. The stakes are substantial: data quality issues represent one of the most common causes of regulatory findings, potentially delaying or preventing drug approval.
Traditional approaches to clinical data cleaning rely heavily on manual processes that have remained fundamentally unchanged for decades. Clinical data managers deploy programmatic edit checks to identify basic inconsistencies, while medical monitors conduct labor-intensive reviews of individual patient narratives, laboratory values, and case report forms (CRFs). This manual approach suffers from several critical limitations: inconsistent application of clinical judgment across reviewers, inability to scale with increasing data volumes, and susceptibility to human error during repetitive tasks. Furthermore, the cognitive burden of synthesizing information across multiple data sources often leads to missed relationships and delayed identification of safety signals.
@@
Despite this promise, the integration of AI into clinical workflows faces significant challenges. Regulatory bodies require transparent, auditable processes for safety-critical decisions. Clinical reviewers express concerns about ”black-box” algorithms making medical judgments. Perhaps most importantly, any AI system must demonstrate not only efficiency gains, but also maintenance or improvement of data quality standards to gain acceptance in this highly regulated environment. Here, we present Octozi, an AI-assisted clinical data cleaning platform designed to address these challenges through a hybrid approach combining machine learning with expert-encoded clinical logic. Our system leverages unified data pipelines to continuously ingest and contextualize diverse clinical trial data, presenting AI-generated insights through an intuitive interface that maintains human oversight while dramatically improving efficiency. We hypothesized that this human-AI collaboration model would improve both the speed and accuracy of clinical data review compared to traditional methods.
@@
To rigorously test this hypothesis, we conducted a controlled experimental study with experienced clinical reviewers, comparing the performance between traditional spreadsheet-based review and AI-assisted workflows. Our results demonstrate that AI assistance fundamentally can shift the focus of clinical development teams from reactive data cleaning to proactive safety and efficacy monitoring. By dramatically reducing the time spent on routine data reconciliation, AI enables teams to engage in higher value activities: identifying safety signals earlier, optimizing trial protocols in real time, and making more informed decisions about patient care. This transformation not only accelerates database-lock but enables more adaptive, responsive trials that can pivot based on emerging data patterns, ultimately reducing both development risk and time to market for life-saving therapies.
@@ 2.1. Study Design and Statistical Considerations
We designed a within-subjects controlled experiment to evaluate the impact of AI assistance on clinical data review performance. This design allowed each participant to serve as their own control, minimizing inter-individual variability while maximizing statistical power. Prior to the study, we conducted an a priori power analysis to determine an appropriate sample size for detecting a meaningful improvement in clinical data cleaning throughput. Assuming a conservative large effect size (Cohen’s d = 1.0), a two-sided paired t-test with $\alpha$ = 0.05 would require a minimum of 10 participants to achieve 80% statistical power. This threshold was selected to detect at least a 50% improvement in throughput, which we considered a clinically meaningful gain based on early pilot observations.
@@
Study participants were recruited from a diverse pool of clinical research professionals through targeted outreach to pharmaceutical companies, contract research organizations, and academic medical centers. Eligibility criteria included: (1) minimum two years of experience in clinical data review roles, (2) proficiency in adverse event adjudication, and (3) familiarity with oncology clinical data cleaning. Pre-study assessments captured detailed professional backgrounds including clinical data cleaning experience, familiarity with specific data management systems, and work experience.
@@ 2.3. Experimental Protocol
The study protocol was designed to minimize bias and consisted of three carefully structured phases. The baseline assessment phase began with participants receiving standardized training on the dataset structure and common discrepancy types. Following a 10-minute briefing period, participants conducted manual reviews using industry-standard spreadsheet tools, with time tracking initiated immediately after the orientation and participant setup of their spreadsheet tool environment. This approach replicated typical clinical data management workflows used across the pharmaceutical industry.
@@
The final phase consisted of standardized assessment using two validated instruments. The NASA Task Load Index (NASA-TLX) was administered to quantify the cognitive workload associated with each review method across six dimensions: mental demand, physical demand, temporal demand, performance, effort, and frustration. This multidimensional assessment was particularly relevant given the cognitive intensity of clinical data review and the potential for AI to reduce mental burden. The System Usability Scale (SUS) provided a standardized measure of the Octozi platform’s usability, generating a composite score that enabled benchmarking against industry standards for enterprise software. Together, these instruments captured both the cognitive impact of AI assistance and the practical usability of the technology, providing crucial insights into the feasibility of widespread adoption in clinical research settings.
@@
Figure 1. Individual participant throughput in correctly cleaned clinical data listings using traditional spreadsheets versus the AI-assisted Octozi platform. Across all participants, Octozi substantially increased data cleaning throughput, with a mean improvement of 6.03-fold (range: 3.7x to 15x). Participants cleaned a mean of 3.4 data points (range: 1–6) with spreadsheets and a median of 20.5 data points (range: 7–35) using Octozi within the same 30-minute review period. Improvements were observed consistently across all participants, highlighting the effectiveness of AI assistance in clinical data review tasks.
2.4. Discrepancy Detection Algorithms
@@
Octozi employs proprietary large language models (LLMs) derived from Llama 4, specifically fine-tuned for clinical trial data cleaning tasks. The discrepancy detection engine combines these LLMs with a suite of heuristic algorithms, systematically developed in collaboration with expert clinical data reviewers. The hybrid approach enables the platform to efficiently identify complex data inconsistencies by integrating domain-specific clinical reasoning with advanced language model capabilities. When evaluated on the annotated synthetic dataset, the discrepancy detection algorithms achieved a classification accuracy of 83.6%, with a recall of 97.5% and precision of 77.2%. Importantly, the accuracy metric reflects correct prediction of the specific discrepancy category, rather than simple anomaly detection, underscoring the model’s capacity to replicate nuanced clinical decision-making patterns.
@@ 3.1. Data Cleaning Efficiency
Notably, the visualization of participant trajectories (Fig. 2) demonstrates that every single participant experienced both improved accuracy and increased throughput when using Octozi, with no instances of speed-accuracy tradeoffs. The consistent upward and rightward movement of all participant data points illustrates the dual benefit of AI assistance, faster and more accurate data cleaning. The worst performer using Octozi performed better than the best participant using spreadsheets, highlighting the substantial and consistent improvement.
@@
3.2. Doubling of Correct Discrepancy Accuracy
The confusion matrices reveal substantial improvements to clinical data cleaning in terms of classification accuracy (Fig. 3). Overall accuracy improved from 45.3% with manual review to 91.5% with Octozi assistance, representing a 46.2 percentage point absolute improvement or a 2.02-fold relative improvement. This enhancement was driven by dramatic improvements in precision, which increased from 44.6% to 93.2% (48.6 percentage point improvement), indicating that AI assistance nearly eliminated false positive identifications.
The F1-score, which balances precision and recall, improved from 58.5 to 91.0, demonstrating that AI assistance enhanced both the completeness and accuracy of discrepancy detection. Interestingly, recall remained high in both conditions (85.2% vs 88.9%), suggesting that experienced reviewers were already adept at recognizing potential issues, but AI assistance helped them correctly distinguish true discrepancies from false alarms.
@@ 3.4. Statistical Significance and Effect Sizes
All primary performance differences between methods achieved statistical significance with large effect sizes. The accuracy improvement yielded a Cohen’s d of 2.8, indicating an extremely large effect size that represents meaningful practical significance beyond statistical significance. Similarly, throughput improvements demonstrated effect sizes exceeding 3.0, confirming that observed differences reflect genuine performance advantages rather than statistical artifacts. Paired t-test analysis confirmed that improvements were consistent across participants (p ¡ 0.001 for all primary measures), with 95% confidence intervals that excluded the possibility of equivalent performance between methods. The magnitude and consistency of observed improvements provide strong evidence for the superiority of AI-assisted clinical data review in both efficiency and accuracy domains.

### v2

# Leveraging AI to Accelerate Medical Data Cleaning: A Comparative Study of AI-Assisted vs. Traditional Methods

@@ Abstract.
Clinical trial data cleaning represents a critical bottleneck in drug development, with manual review processes struggling to manage exponentially increasing data volumes and complexity. This paper presents Octozi, an artificial intelligence-assisted platform that combines large language models with domain-specific heuristics to transform medical data review. In a controlled experimental study with experienced medical reviewers (n=10), we demonstrate that AI assistance increased data cleaning throughput by 6.03-fold while simultaneously decreasing cleaning errors from 54.67% to 8.48% (a 6.44-fold improvement). Crucially, the system reduced false positive queries by 15.48-fold, minimizing unnecessary site burden. Economic analysis of a representative Phase III oncology trial reveals potential cost savings of \$5.1 million, primarily driven by accelerated database lock timelines (5-day reduction saving \$4.4M), improved medical review efficiency (\$420K savings), and reduced query management burden (\$288K savings). These improvements were consistent across reviewers regardless of experience level, suggesting broad applicability. Our findings indicate that AI-assisted approaches can address fundamental inefficiencies in clinical trial operations, potentially accelerating drug development timelines such as database lock by 33% while maintaining regulatory compliance and significantly reducing operational costs. This work establishes a framework for integrating AI into safety-critical clinical workflows and demonstrates the transformative potential of human-AI collaboration in pharmaceutical clinical trials.
@@ 1. Introduction
The pharmaceutical industry faces an unprecedented data management crisis. Clinical trials continue to generate more data than a decade ago (Taves, 2010), yet the fundamental processes for ensuring data quality remain largely unchanged since the advent of electronic data capture (EDC). This disconnect between data volume and processing capability creates substantial delays in drug development, with each day of database lock delay costing \$840,000 in lost revenue across an oncology product’s lifecycle (Smith et al., 2024). More critically, inadequate data cleaning risks missing crucial safety signals that could impact patient welfare.
Medical data cleaning, the systematic process of reviewing, validating, and reconciling data collected from investigative sites, serves as the foundation for regulatory submissions to authorities including the US Food and Drug Administration (FDA) and European Medicines Agency (EMA). This process extends beyond simple error detection to encompass complex medical judgment, requiring specialized expertise to identify clinically meaningful discrepancies while maintaining regulatory compliance. The stakes are substantial: data quality issues represent one of the most common causes of regulatory findings, potentially delaying or preventing drug approval.
Traditional approaches to medical data cleaning rely heavily on manual processes that have remained fundamentally unchanged for decades. Medical data managers deploy programmatic edit checks to identify basic inconsistencies, while medical monitors conduct labor-intensive reviews of individual patient narratives, laboratory values, and case report forms (CRFs). This manual approach suffers from several critical limitations: inconsistent application of clinical judgment across reviewers, inability to scale with increasing data volumes, and susceptibility to human error during repetitive tasks. Furthermore, the cognitive burden of synthesizing information across multiple data sources often leads to missed relationships and delayed identification of safety signals.
@@
Despite this promise, the integration of AI into clinical workflows faces significant challenges. Regulatory bodies require transparent, auditable processes for safety-critical decisions. Medical reviewers express concerns about ”black-box” algorithms making medical judgments. Perhaps most importantly, any AI system must demonstrate not only efficiency gains, but also maintenance or improvement of data quality standards to gain acceptance in this highly regulated environment. Here, we present Octozi, an AI-assisted medical data cleaning platform designed to address these challenges through a hybrid approach combining machine learning with expert-encoded clinical logic. Our system leverages unified data pipelines to continuously ingest and contextualize diverse clinical trial data, presenting AI-generated insights through an intuitive interface that maintains human oversight while dramatically improving efficiency. We hypothesized that this human-AI collaboration model would improve both the speed and accuracy of medical data review compared to traditional methods.
@@
To rigorously test this hypothesis, we conducted a controlled experimental study with experienced medical reviewers, comparing the performance between traditional spreadsheet-based review and AI-assisted workflows. Our results demonstrate that AI assistance fundamentally can shift the focus of clinical development teams from reactive data cleaning to proactive safety and efficacy monitoring. By dramatically reducing the time spent on routine data reconciliation, AI enables teams to engage in higher value activities: identifying safety signals earlier, optimizing trial protocols in real time, and making more informed decisions about patient care. This transformation not only accelerates database-lock but enables more adaptive, responsive trials that can pivot based on emerging data patterns, ultimately reducing both development risk and time to market for life-saving therapies.
2. Related Work
Clinical trial data quality remains a persistent challenge as data volumes and complexity continue to grow exponentially. Modern oncology trials generate millions of data points across dozens of interconnected systems, yet the fundamental processes for ensuring data quality rely on methods developed decades ago for much smaller datasets. This mismatch between data scale and review capabilities creates substantial delays in drug development and risks missing critical safety signals. Various technological approaches have emerged to modernize data review, though each faces significant limitations that prevent comprehensive adoption.
Visualization platforms like JReview and Spotfire enable identification of numerical outliers through scatter plots and heat maps, but face fundamental constraints. These tools cannot process unstructured text data, which comprises the majority of critical data fields such as adverse event narratives and concomitant medication descriptions. Moreover, statistical outliers identified through visualization do not necessarily represent clinically meaningful discrepancies. A laboratory value may be statistically unusual but clinically appropriate given a patient’s medical context, while dangerous drug interactions may appear normal in aggregate visualizations.
Recent advances in machine learning have enabled more sophisticated programmatic edit checks that use historical data to calibrate thresholds and reduce false positives. However, these systems remain limited by Boolean logic that cannot extend to unstructured text. They cannot perform the complex medical reasoning required to determine whether an adverse event is consistent with a patient’s medical history. The inability to process narrative descriptions means these systems miss critical relationships that are often documented only in free-text fields.
Risk-Based Quality Management (RBQM) approaches, endorsed by FDA and EMA in 2013, use statistical algorithms to identify sites or patterns deviating from expected norms (Stoltz, 2013). While conceptually appealing, these methods require substantial sample sizes to achieve adequate statistical power, rendering them less effective for early trial phases or rare disease studies. Furthermore, like visualization approaches, statistical monitoring cannot process narrative text, and statistical significance does not equate to clinical significance. An RBQM solution, for instance, could not alert a medical reviewer that a concomitant medication was inappropriate for treating a documented adverse event, as this requires understanding the medical relationship between free-text descriptions rather than detecting statistical anomalies.
Our approach with Octozi represents a fundamental departure from these existing paradigms. Rather than adapting general-purpose tools to clinical data, we developed a purpose-built system that combines large language models with deterministic clinical algorithms and external knowledge sources such as FDA drug labels. This hybrid architecture enables several key innovations. First, our LLM-based approach natively processes both structured and unstructured data, identifying discrepancies invisible to traditional methods. For example, recognizing when a narrative description of ”severe nausea requiring hospitalization” contradicts a coded severity of ”mild.” Second, integration of external medical knowledge enables true clinical reasoning rather than pattern matching, identifying contraindications and drug interactions that would appear normal in statistical analyses. Third, unlike statistical methods requiring minimum sample sizes, our system scales seamlessly from single-patient review to population analysis through consistent application of clinical logic. Finally, the platform contextualizes data within the patient’s journey, differentiating between expected clinical progressions and true discrepancies recognizing, for example, that platelet drops after the initiation of chemotherapy are expected while the same change without clinical context warrants review. This purpose-built approach addresses the fundamental mismatch between existing tools and medical review requirements, offering a solution that enhances clinical expertise while dramatically improving efficiency and consistency.
@@ 3.1. Study Design and Statistical Considerations
We designed a within-subjects controlled experiment to evaluate the impact of AI assistance on medical data review performance. This design allowed each participant to serve as their own control, minimizing inter-individual variability and thereby ensuring that observed differences reflected method effects rather than participant differences. Prior to the study, we conducted an a priori power analysis to determine an appropriate sample size for detecting a meaningful improvement in clinical data cleaning throughput. Assuming a conservative large effect size (Cohen’s d = 1.0), a two-sided paired t-test with $\alpha$ = 0.05 would require a minimum of 10 participants to achieve 80% statistical power. This threshold was selected to detect at least a 50% improvement in throughput, which we considered a clinically meaningful gain based on early pilot observations.
@@
Study participants were recruited from a diverse pool of clinical research professionals through targeted outreach to pharmaceutical companies, contract research organizations, and academic medical centers. Eligibility criteria included: (1) a minimum two years of experience in medical data review roles, (2) proficiency in adverse event adjudication, and (3) familiarity with oncology medical data cleaning. Pre-study assessments (Appendix B) captured detailed professional backgrounds including medical data cleaning experience, familiarity with specific data management systems, and work experience.
@@ 3.3. Experimental Protocol
The study protocol was designed to minimize bias and consisted of three carefully structured phases. The baseline assessment phase began with participants receiving standardized training on the dataset structure and common discrepancy types. Following a 10-minute briefing period, participants conducted manual reviews using industry-standard spreadsheet tools, with time tracking initiated immediately after the orientation and participant setup of their spreadsheet tool environment. This approach replicated typical medical data review workflows used across the pharmaceutical industry.
@@
The final phase consisted of standardized assessment using two validated surveys. The inverse NASA Task Load Index (NASA-TLX) was administered to quantify the cognitive workload associated with each review method across six dimensions: mental demand, physical demand, temporal demand, performance, effort, and frustration. This multidimensional assessment was particularly relevant given the cognitive intensity of medical data review and the potential for AI to reduce mental burden. The System Usability Scale (SUS) provided a standardized measure of the Octozi platform’s usability, generating a composite score that enabled benchmarking against industry standards for enterprise software. Together, these instruments captured both the cognitive impact of AI assistance and the practical usability of the technology, providing crucial insights into the feasibility of widespread adoption in clinical research settings.
@@
Figure 1. Individual participant throughput in correctly cleaned clinical data listings using traditional spreadsheets versus the AI-assisted Octozi platform. Across all participants, Octozi substantially increased data cleaning throughput, with a mean improvement of 6.03-fold (range: 3.7x to 15x). Participants cleaned a mean of 3.4 data points (range: 1–6) with spreadsheets and a median of 20.5 data points (range: 7–35) using Octozi within the same 30-minute review period. Improvements were observed consistently across all participants, highlighting the effectiveness of AI assistance in medical data review tasks.
3.4. AI-Assisted Medical Data Review Platform
The Octozi platform represents a comprehensive AI-assisted solution designed to transform medical data review through intelligent data integration, contextual analysis, and automated discrepancy detection. The system architecture comprises four core components that work synergistically to enhance reviewer efficiency and accuracy.
3.4.1. Data Integration and Harmonization
The platform ingests and harmonizes data from multiple clinical trial sources, including electronic data capture (EDC) systems, safety databases, central laboratory systems, and patient-reported outcome (PRO) platforms. Unlike traditional approaches that require reviewers to navigate between disparate systems, Octozi creates a unified data model that preserves source traceability while enabling cross-domain analysis.
3.4.2. Contextual Intelligence Engine
A distinguishing feature of the platform is its ability to automatically contextualize each data point within the patient’s clinical journey. The system employs temporal reasoning algorithms to identify clinically relevant relationships between events, such as:
  • Temporal associations between laboratory abnormalities and medication changes (e.g., elevated liver enzymes occurring 72 hours after initiating a hepatotoxic medication)
  • Pattern recognition across multiple parameters (e.g., correlating weight gain, edema reports, and elevated BNP levels to identify potential heart failure)
This contextual analysis extends beyond simple temporal proximity to incorporate clinical significance scoring based on established medical guidelines and expert-encoded rules. For example, an abnormal hemoglobin level in a patient with a history of anemia is identified with a different clinical significance than a patient experiencing a sudden drop in hemoglobin.
3.4.3. Intelligent Discrepancy Detection
@@
Octozi employs proprietary large language models (LLMs) derived from Llama 4, specifically fine-tuned for clinical trial data cleaning tasks (met, [n. d.]). The discrepancy detection engine combines these LLMs with a suite of heuristic algorithms, systematically developed in collaboration with expert medical data reviewers. The hybrid approach enables the platform to efficiently identify complex data inconsistencies by integrating domain-specific clinical reasoning with advanced language model capabilities. When evaluated on the annotated synthetic dataset, the discrepancy detection algorithms achieved a classification accuracy of 83.6%, with a recall of 97.5% and precision of 77.2%. Importantly, the accuracy metric reflects the correct prediction of the specific category of discrepancies, rather than the simple detection of anomalies, underscoring the model’s ability to replicate nuanced clinical decision-making patterns.
3.4.4. Interactive Review Interface
The user interface presents identified discrepancies and contextual information through multiple complementary views. A tabular view provides structured access to raw data with color-coded severity indicators and filtering capabilities. Moreover, a narrative view generates human-readable summaries that explain the clinical context and rationale for the flagged items, mimicking the mental model of experienced reviewers. This dual presentation format accommodates different cognitive styles and review scenarios.
The platform also incorporates query generation assistance, providing suggested query text based on the specific discrepancy type and regulatory requirements. This feature standardizes query language across reviewers while reducing the time required for query composition. Reviewers retain full control to modify or reject suggested queries, maintaining human oversight of all decisions.
Visual examples of the platform’s interface and workflow are provided in Appendix B, demonstrating how the system presents complex clinical relationships in an intuitive, actionable format that accelerates decision-making while maintaining audit trails for regulatory compliance.
@@ 4.1. Data Cleaning Efficiency
Notably, the visualization of participant trajectories (Fig. 2) demonstrates that every single participant experienced both improved accuracy and increased throughput when using Octozi, with no observed traded-off between speed and accuracy. The consistent upward and rightward movement of all participant data points illustrates the dual benefit of AI assistance, faster and more accurate data cleaning. The worst performer using Octozi performed better than the best participant using spreadsheets, highlighting the substantial and consistent improvement.
@@
4.2. Reduction in Medical Review Error Rates
The confusion matrices reveal a transformative improvement in clinical data cleaning accuracy through AI assistance (Fig. 4). Most notably, the error rate plummeted from 54.7% with manual review to just 8.5% with Octozi assistance, representing a 6.44-fold reduction in errors. This dramatic improvement means that reviewers using AI assistance made mistakes on fewer than 1 in 11 data points, compared to errors on more than half of all data points when using traditional spreadsheet methods.
This error reduction was driven by substantial improvements across all performance metrics. Overall accuracy improved from 45.3% to 91.5%, representing a 46.2 percentage point absolute improvement. The enhancement was particularly pronounced in precision, which increased from 44.6% to 93.2% (48.6 percentage point improvement), indicating that AI assistance nearly eliminated false positive identifications. When reviewers flagged a discrepancy using Octozi, they were correct 93% of the time, compared to being wrong more often than right (44.6% precision) with manual methods.
The F1-score, which balances precision and recall, improved from 58.5 to 91.0 (32.4 point increase), demonstrating that AI assistance enhanced both the completeness and accuracy of discrepancy detection. Interestingly, recall remained high in both conditions (85.2% vs 88.9%), suggesting that experienced reviewers were already adept at recognizing potential issues, but AI assistance helped them correctly distinguish true discrepancies from false alarms. The 6.44-fold reduction in error rate represents not just an incremental improvement but a fundamental shift in review reliability, transforming clinical data cleaning from an error-prone process to a highly accurate operation.
@@ 4.4. Statistical Significance and Effect Sizes
All primary performance differences between methods achieved statistical significance with large effect sizes. The accuracy improvement yielded a Cohen’s d of 2.8, indicating an extremely large effect size that represents meaningful practical significance beyond statistical significance. Similarly, throughput improvements demonstrated effect sizes exceeding 3.0, confirming that observed differences reflect genuine performance advantages rather than statistical artifacts. Paired t-test analysis confirmed that improvements were consistent across participants (p ¡ 0.001 for all primary measures), with 95% confidence intervals that excluded the possibility of equivalent performance between methods. The magnitude and consistency of observed improvements provide strong evidence for the superiority of AI-assisted medical data review in both efficiency and accuracy domains.
Figure 3. System Usability Scale (SUS) scores demonstrate superior usability of Octozi (88.3) compared to traditional spreadsheets (55.4), with scores above 68 considered acceptable and above 80 considered excellent. Inverse NASA Task Load Index (NASA-TLX) scores reveal a 3.2-fold reduction in cognitive workload with Octozi (83.3) versus spreadsheets (25.7), where higher scores indicate less mental demand, physical demand, temporal demand, performance concerns, effort, and frustration.